OpenStudy (anonymous):

Algebra 2, logarithms??

OpenStudy (anonymous):

If \[f(x)=\log_2 (x+4)\] what is\[f ^{-1}(3)\] a. 0 b. 2 c. 4 d. 8

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

Replace f(x) with 3 and solve for x.

OpenStudy (mathstudent55):

\(3 = \log_2 (x + 4) \) Now use the definition of log.

OpenStudy (anonymous):

x=4

OpenStudy (mathstudent55):

\(2 ^ 3 = x + 4\) \(8 = x + 4\) \(x = 4\) You are correct.

OpenStudy (anonymous):

I have a question on parent functions. I've worked it out and somehow I ended up with g(x)-f(x)=all real numbers.... But that doesn't make any sense at all